OpenStudy (anonymous):

what is LXIX in roman numerals HELP !!!!! I REALLY NEED HELP and PLEASE EXPLAiN HOW U GET THE ANSWER

OpenStudy (anonymous):

hlelp plzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzz

OpenStudy (anonymous):

LXIX in Roman Numerals is 69. L + X + IX = 50 + 10 +9 = 69 L=50 X=10 IX = 9 (one number before ten) Hope that helps.